# Understanding Outer Optimizers in Local SGD: Learning Rates, Momentum, and Acceleration

### v1

@@ 3.2 Main convergence result
Theorem 3.5.
Suppose that Assumptions˜3.1 and 3.2 hold. Suppose that the outer update is gradient descent with momentum, $\mathrm{OuterUpdate}(x_{r},-\Delta_{r,H})=x_{r}+\gamma\Delta_{r,H}+\mu(x_{r}-x_{r-1})$ with momentum parameter $\mu\in[0,1)$ and the local update is gradient descent $\mathrm{LocalUpdate}(y,g)=y-\eta g$ in Algorithm˜1. Let the step sizes $\eta,\gamma$ satisfy $\eta L\left(1+\left(\frac{\gamma}{1-\mu}-1\right)_{+}H\right)\leq\frac{1}{4}$ and $\frac{\eta\gamma\mu LH}{1-\mu}\leq\frac{1}{16}$. Then after $R$ rounds of communication, the averaged iterate satisfies
$\begin{split}\mathbb{E}\left[f(\overline{y})\right]-f(x_{*})&\leq\mathcal{O}\left(\frac{(1-\mu){\left\lVert z_{0}-x_{\ast}\right\rVert}^{2}}{\eta\gamma HR}+L\eta^{2}\sigma^{2}H+\frac{\eta\sigma^{2}\left(1+(\frac{\gamma}{1-\mu}-1)_{+}\right)}{M}+\frac{\eta\gamma\mu}{1-\mu}\frac{\sigma^{2}}{M}\right),\end{split}$
@@
The proof is provided in Section˜B.3. Theorem 3.5 shows the requirement on the outer stepsize is relaxed from a requirement on $\gamma$ to a requirement on $\frac{\gamma}{1-\mu}$, allowing us to reap the same benefits of $\gamma>1$ observed earlier if we also tune $\mu$. This benefit was first observed in [53] for nonconvex optimization with small local stepsize $\eta$ provided we use an additional momentum buffer. Our work gives direct theoretical support to this observation even with a single momentum buffer and allowing for large $\eta$.
@@ 3.3 Convergence with accelerated outer optimizer
We now consider the use of acceleration. To the best of our knowledge, the combination of an accelerated outer optimizer with an unaccelerated inner optimizer, as in e.g. DiLoCo [10, 23], has not been analyzed in the literature before. We take steps towards addressing this gap and understanding the convergence properties of such algorithms by considering Nesterov’s accelerated gradient descent [43] as the outer optimizer and (stochastic) gradient descent as the inner optimizer. The following theorem gives a convergence guarantee for this setting.
Theorem 3.6.
Suppose that Assumptions˜3.1 and 3.2 hold and the stepsizes satisfy $2L\eta\leq 1$ and $\gamma\leq 1$. Suppose that the outer update is accelerated gradient descent with Nesterov momentum as follows
$u_{r+1}=x_{r}-\eta\Delta_{r,H},$ $z_{r+1}=z_{r}-\gamma_{r}\eta\Delta_{r,H},$ $x_{r+1}=(1-\tau_{r+1})u_{r+1}+\tau_{r+1}z_{r+1},$
@@
Corollary 3.7.
In the same setting as Theorem˜3.6, setting $\gamma=1$ in Equation˜7, and choosing
$\eta=\min\left\{\frac{1}{2L},\left(\frac{2MD^{2}}{R^{3}L\sigma^{2}H^{2}}\right)^{1/3},\left(\frac{4D^{2}}{3R^{3}L^{2}\sigma^{2}H^{3}}\right)^{1/4},\sqrt{\frac{4MD^{2}}{R^{3}H\sigma^{2}}}\right\},$
@@
$\mathbb{E}[f(u_{R})]-f(x_{*})$ $\leq\mathcal{O}\Bigl{(}\frac{LD^{2}}{R^{2}H}+\frac{L^{1/3}\sigma^{2/3}D^{4/3}}{RM^{1/3}H^{1/3}}+\frac{L^{1/2}\sigma^{1/2}D^{3/2}}{R^{5/4}H^{1/4}}+\frac{\sigma D}{\sqrt{MRH}}\Bigr{)}.$ (8)
Equation˜8 shows that in the absence of noise, we obtain a rate accelerated in $R$ but not $H$. This intuitively makes sense, since we do acceleration only in the outer loop. In the presence of noise, we have in the worst-case the unimprovable $\frac{\sigma D}{\sqrt{MRH}}$ term and two additional noise terms that characterize the drift suffered by this algorithm. Notably, the drift terms have much better dependence on $R$ compared to Local SGD, as given by Equation˜2. [60] analyze FedAC, an accelerated variant of Local SGD that uses acceleration locally and applies simple averaging as the outer optimizer. Their algorithm enjoys the convergence rate
$\mathbb{E}[f(x_{\mathrm{out}})]-f(x_{\ast})\leq\mathcal{O}\left(\frac{LD^{2}}{R^{2}H}+\frac{L^{1/3}\sigma^{2/3}D^{4/3}}{RH^{1/3}}+\frac{L^{1/2}\sigma^{1/2}D^{3/2}}{RH^{1/4}}+\frac{\sigma D}{\sqrt{MRH}}\right).$
@@ 3.4 Data-dependent convergence result
To further understand the role of the outer stepsize, we now present a data-dependent, high-probability guarantee for Generalized Local SGD in Theorem 3.8, compared to the rather worst-case analysis of Theorem 3.3. This analysis may also provide insights into practical tuning of the outer learning rate
Theorem 3.8.
Suppose that Assumptions˜3.1 and 3.2 hold. Then in Algorithm 1 with outer update $x=x-\gamma\Delta$ and local update $y=y-\eta g$, if the local stepsize satisfies $\eta\leq\frac{1}{L}$ then with probability at least $1-\delta$ the iterates generated satisfy
$\begin{split}&f\left(\frac{1}{RH}\sum_{r=0}^{R-1}\sum_{h=0}^{H-1}y_{r,h}\right)-f(x_{\ast})\leq\tilde{\mathcal{O}}\Biggl{(}\frac{{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\gamma\eta RH}+\frac{\gamma\eta}{RH}\sum_{r,h}{\left\lVert g_{r,h}\right\rVert}^{2}+\gamma\eta\sigma^{2}\\ &+\frac{\left\lvert 1-\gamma\right\rvert\eta}{RH}\sum_{r}\left(\sum_{h}\left\lVert g_{r,h}\right\rVert\right)^{2}+\frac{\eta}{\gamma H}\left(\frac{1}{M}\max_{r}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}+\eta\sigma\sqrt{\frac{1}{MR}\sum_{m,r,h}{\left\lVert g_{m,r,h}\right\rVert}^{2}}\Biggr{)}.\end{split}$
The proof of Theorem 3.8 is provided in Section˜B.5. Compared to Theorem 3.3, the guarantee we obtain here is weaker in some areas, e.g., the variance term $\gamma\eta\sigma^{2}$ does not benefit from increasing $M$. On the other hand, this guarantee is a high-probability and data-dependent guarantee. To the best of our knowledge, this is the first high-probability convergence guarantee for Local SGD in the literature. Theorem 3.8 allows us to observe another potential benefit of using $\gamma\neq 1$. To see how, let us make the simplifying assumption that $\left\lVert\hat{g}_{r,h}\right\rVert\approxeq G_{1}$ and $\left\lVert g_{m,r,h}\right\rVert\approxeq G_{2}$. Observe that by the triangle inequality we have $G_{1}\leq G_{2}$, but in fact $G_{1}$ can be significantly smaller than $G_{2}$, particularly in the later stages of the optimization process, due to the variance reduction effect of averaging together the gradients on different nodes. Then we can rewrite the above guarantee as
@@ 4 Experiments
Figure 1: Effect of varying $\sigma$ and $\gamma$ for quadratic problem
@@ 4.2 Transformer pretraining
Setup Following the DiLoCo paper [10], we experiment using a Chinchilla decoder transformer [18] on the C4 dataset [47]. The architecture hyperparameters are identical from the DiLoCo paper [10] and are given in Section˜A.1.1. We fix the batch size at $512$ and the sequence length at $1024$. We experiment at different scales, from 150 million to 1 billion parameters. For all experiments, the inner optimizer is AdamW [36] trained with a cosine learning rate schedule defined across the total amount of steps. The inner optimizer state is never shared across replicas, and is passed from one round to the other.
@@
Figure 2: Scaling distributed pretraining, at 150M, 400M, and 1B parameters.
Methods We compare three distributed methods, using different outer optimizers: SGD(lr=1) (equivalent to simple averaging of local models [39]), Nesterov (equivalent to DiLoCo [10]), and ScheduleFree-SGD (SF-SGD) [8]. We use SF-SGD to substitute for outer learning rate scheduling, though it still requires tuning hyperparameters. We also include two “high-communication" data-parallel baselines: one with the global batch size as the local per-replica batch size used by the distributed methods, and one with the same batch size as the global batch size ($M\times$ the local per-replica batch size) used by the distributed methods. The latter requires either more GPUs and more thus communication, or gradient accumulation and thus more time. The latter also has an equal flops budget as the distributed methods. We tuned all our optimizers on the pretraining setting on a separate validation set . We also considered using SF-Nesterov, but it was hard to tune and unstable.
@@
Table 1: Optimizer hyperparameters for the three evaluated sizes. All are based on the transformer architecture, chinchilla-style [18].
Results Table˜1 gives the optimal hyperparameters per scale, and Figure˜2 gives the perplexity curves. Consistent with the predictions of our theory, we found that an outer learning rate greater than $1.0$ performed best for SF-SGD and a relatively large effective outer learning rate also performed best for Nesterov; Moreover, acceleration consistently improved performance relative to the baseline Local SGD. In the supplementary material, we report the effect of varying the number of local steps (Section˜A.1.2), the number of clients/replicas and different ways of FLOPs allocation (Section˜A.1.3), and gradient variance (Section˜A.1.6).
@@ A.1 Language model pretraining
We study the impact of using various outer optimizers on large language model pretraining. We utilized Chinchilla-style decoder transformer architectures [18] trained on the C4 dataset [47], consistent with common practices in large-scale model training [10]. The following subsections detail the specific hyperparameters, variations in training configurations (such as the number of inner steps and replicas/clients), and analyses of optimizer behavior, including learning rate scheduling and observed gradient cosine similarities.
@@ A.1.1 Hyperparameters details
We show in Table 1 the hyperparameters considered and kept, and in Table 2 the architectural hyperparameters. We tuned all our optimizers on a separate validation set. We also considered using the Schedule-Free Optimizer with Nesterov acceleration on top but it was hard to tune and unstable.
Table 2: Model Configuration for the three evaluated sizes. All are based on the transformer architecture, chinchilla-style [18].
@@ A.1.4 Schedule-free but not tuning-free
The schedule-free method of [8] enables not doing any learning rate scheduling, greatly simplifying training configuration. However, it doesn’t mean it is hyperparameters-tuning-free. Indeed, we found out that we had to extensively tune the initial learning rate (to $2.0$), remove learning rate warm-up contrarily to what is advised, and use a particularly low $b1$ decay: $0.2$, as illustrated in Figure 7.
@@ A.1.5 Pretraining: outer learning rate scheduling
Schedule-free SGD enables not having to manually scheduling the outer learning rate. Therefore, we wondered if we could improve the SotA federated learning baseline, DiLoCo (Nesterov outer optimizer), with an outer learning rate schedule. We investigate in Figure 7 three schedules: constant as in [10], cosine decay, and linear after a plateau. For the latter we consider a constant plateau for 10% and 25% of the total steps. For each method, we also tuned the peak outer learning rate. We don’t use any warm-up in the outer optimization as we always found it to be harmful.
We find that constant outer learning rate is the best performing schedule. It’s unclear how the other schedules are interacting with the inner learning rate scheduling. A possible solution, not investigated in this report, would be to increase the number of inner steps $H$ as the inner learning rate decreases [14].

### v2

@@ 3.2 Main convergence result
Theorem 2.
Suppose that Assumptions˜1 and 2 hold. Suppose that the outer update is gradient descent with momentum, $\mathrm{OuterUpdate}(x_{r},-\Delta_{r,H})=x_{r}+\gamma\Delta_{r,H}+\mu(x_{r}-x_{r-1})$ with momentum parameter $\mu\in[0,1)$ and the local update is gradient descent $\mathrm{LocalUpdate}(y,g)=y-\eta g$ in Algorithm˜1. Let the step sizes $\eta,\gamma$ satisfy $\eta L\left(1+\left(\frac{\gamma}{1-\mu}-1\right)_{+}H\right)\leq\frac{1}{4}$ and $\frac{\eta\gamma\mu LH}{1-\mu}\leq\frac{1}{16}$. Then after $R$ rounds of communication, the averaged iterate satisfies
$\begin{split}\textstyle\mathbb{E}\left[f(\overline{y})\right]-f(x_{*})&\leq\mathcal{O}\left(\frac{(1-\mu){\left\lVert z_{0}-x_{\ast}\right\rVert}^{2}}{\eta\gamma HR}+L\eta^{2}\sigma^{2}H+\frac{\eta\sigma^{2}\max(\frac{\gamma}{1-\mu},1)}{M}+\frac{\eta\gamma\mu}{1-\mu}\frac{\sigma^{2}}{M}\right),\end{split}$
@@
The proof is provided in Section˜B.3. Theorem 2 shows the requirement on the outer stepsize is relaxed from a requirement on $\gamma$ to a requirement on $\frac{\gamma}{1-\mu}$, allowing us to reap the same benefits of $\gamma>1$ observed earlier if we also tune $\mu$. Momentum thus changes the range of stepsizes allowed but does not fundamentally alter the uter stepsize tradeoffs. This benefit was first observed in [sun23_role_server_momen_feder_learn] for nonconvex optimization with small local stepsize $\eta$ provided we use an additional momentum buffer. Our work gives direct theoretical support to this observation even with a single momentum buffer and allowing for large $\eta$.
@@ 3.3 Convergence with accelerated outer optimizer
We now consider the use of acceleration. To the best of our knowledge, the combination of an accelerated outer optimizer with an unaccelerated inner optimizer, as in e.g. DiLoCo [douillard23_diloc, jaghouar24_opend], has not been analyzed in the literature before. We take steps towards addressing this gap and understanding the convergence properties of such algorithms by considering Nesterov’s accelerated gradient descent [nesterov18_lectures_cvx_opt] as the outer optimizer and (stochastic) gradient descent as the inner optimizer. The following theorem gives a convergence guarantee for this setting.
Theorem 3.
Suppose that Assumptions˜1 and 2 hold and the stepsizes satisfy $2L\eta\leq 1$ and $\gamma\leq 1$. Suppose that the outer update is accelerated gradient descent with Nesterov momentum as follows
$\textstyle u_{r+1}=x_{r}-\eta\Delta_{r,H},$ $z_{r+1}=z_{r}-\gamma_{r}\eta\Delta_{r,H},$ $x_{r+1}=(1-\tau_{r+1})u_{r+1}+\tau_{r+1}z_{r+1},$
@@
Corollary 1.
In the same setting as Theorem˜3, setting $\gamma=1$ in Equation˜7, and choosing
$\textstyle\eta=\min\left\{\frac{1}{2L},\left(\frac{2MD^{2}}{R^{3}L\sigma^{2}H^{2}}\right)^{1/3},\left(\frac{4D^{2}}{3R^{3}L^{2}\sigma^{2}H^{3}}\right)^{1/4},\sqrt{\frac{4MD^{2}}{R^{3}H\sigma^{2}}}\right\},$
@@
$\textstyle\mathbb{E}[f(u_{R})]-f(x_{*})$ $\leq\mathcal{O}\Bigl(\frac{LD^{2}}{R^{2}H}+\frac{L^{1/3}\sigma^{2/3}D^{4/3}}{RM^{1/3}H^{1/3}}+\frac{L^{1/2}\sigma^{1/2}D^{3/2}}{R^{5/4}H^{1/4}}+\frac{\sigma D}{\sqrt{MRH}}\Bigr).$ (8)
Equation˜8 shows that in the absence of noise, we obtain a rate accelerated in $R$ but not $H$. This intuitively makes sense, since we do acceleration only in the outer loop. In the presence of noise, we have in the worst-case the unimprovable $\frac{\sigma D}{\sqrt{MRH}}$ term and two additional noise terms that characterize the drift suffered by this algorithm. Notably, the drift terms have much better dependence on $R$ compared to Local SGD, as given by Equation˜2. [yuan20_feder_accel_stoch_gradien_descen] analyze FedAC, an accelerated variant of Local SGD that uses acceleration locally and applies simple averaging as the outer optimizer. Their algorithm enjoys the convergence rate
$\textstyle\mathbb{E}[f(x_{\mathrm{out}})]-f(x_{\ast})\leq\mathcal{O}\left(\frac{LD^{2}}{R^{2}H}+\frac{L^{1/3}\sigma^{2/3}D^{4/3}}{RH^{1/3}}+\frac{L^{1/2}\sigma^{1/2}D^{3/2}}{RH^{1/4}}+\frac{\sigma D}{\sqrt{MRH}}\right).$
@@ 3.4 Data-dependent convergence result
To further understand the role of the outer stepsize, we now present a data-dependent, high-probability guarantee for Generalized Local SGD in Theorem 4, compared to the rather worst-case analysis of Theorem 1. This analysis may also provide insights into practical tuning of the outer learning rate
Theorem 4.
Suppose that Assumptions˜1 and 2 hold. Then in Algorithm 1 with outer update $x=x-\gamma\Delta$ and local update $y=y-\eta g$, if the local stepsize satisfies $\eta\leq\frac{1}{L}$ then with probability at least $1-\delta$ the iterates generated satisfy
$\textstyle\begin{split}\textstyle&f\left(\frac{1}{RH}\sum_{r=0}^{R-1}\sum_{h=0}^{H-1}y_{r,h}\right)-f(x_{\ast})\leq\tilde{\mathcal{O}}\Biggl(\frac{{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\gamma\eta RH}+\frac{\gamma\eta}{RH}\sum_{r,h}{\left\lVert g_{r,h}\right\rVert}^{2}+\gamma\eta\sigma^{2}\\ &+\frac{\left\lvert 1-\gamma\right\rvert\eta}{RH}\sum_{r}\left(\sum_{h}\left\lVert g_{r,h}\right\rVert\right)^{2}+\frac{\eta}{\gamma H}\left(\frac{1}{M}\max_{r}\sum_{m,h}\left\lVert g_{m,r,h}\right\rVert\right)^{2}+\eta\sigma\sqrt{\frac{1}{MR}\sum_{m,r,h}{\left\lVert g_{m,r,h}\right\rVert}^{2}}\Biggr).\end{split}$
The proof of Theorem 4 is provided in Section˜B.5. Compared to Theorem 1, the guarantee we obtain here is weaker in some areas, e.g., the variance term $\gamma\eta\sigma^{2}$ does not benefit from increasing $M$. On the other hand, this guarantee is a high-probability and data-dependent guarantee. To the best of our knowledge, this is the first high-probability convergence guarantee for Local SGD in the literature. Theorem 4 allows us to observe another potential benefit of using $\gamma\neq 1$. To see how, let us make the simplifying assumption that $\left\lVert\hat{g}_{r,h}\right\rVert\approxeq G_{1}$ and $\left\lVert g_{m,r,h}\right\rVert\approxeq G_{2}$. Observe that by the triangle inequality we have $G_{1}\leq G_{2}$, but in fact $G_{1}$ can be significantly smaller than $G_{2}$, particularly in the later stages of the optimization process, due to the variance reduction effect of averaging together the gradients on different nodes. Then we can rewrite the above guarantee as
@@ 4 Experiments
Figure 1: Effect of varying noise magnitude $\sigma$ and outer learning rate $\gamma$ for quadratic optimization.
@@ 4.2 Transformer pretraining
Setup Following the DiLoCo paper [douillard23_diloc], we experiment using a Chinchilla decoder transformer [hoffmann2022chinchilla] on the C4 dataset [c4]. The architecture hyperparameters are identical from the DiLoCo paper [douillard23_diloc] and are given in Section˜A.1.1. We fix the batch size at $512$ and the sequence length at $1024$. We experiment at different scales, from 150 million to 1 billion parameters. For all experiments, the inner optimizer is AdamW [loshchilov17_decoup_weigh_decay_regul] trained with a cosine learning rate schedule defined across the total amount of steps. The inner optimizer state is never shared across replicas, and is passed from one round to the other.
@@
Figure 2: Scaling distributed pretraining, at 150M, 400M, and 1B parameters. The x-axis shows the total number of training steps, including both local and communication steps. The y-axis shows the perplexity achieved by each method. Legend represents final perplexity values.
Methods We compare three distributed methods, using different outer optimizers: SGD(lr=1) (equivalent to simple averaging of local models [mcmahan2021fedavg]), Nesterov (equivalent to DiLoCo [douillard23_diloc]), and ScheduleFree-SGD (SF-SGD) [defazio24_road_less_sched]. We use SF-SGD to substitute for outer learning rate scheduling, though it still requires tuning hyperparameters. We also include two “high-communication" data-parallel baselines: one with the global batch size as the local per-replica batch size used by the distributed methods, and one with the same batch size as the global batch size ($M\times$ the local per-replica batch size) used by the distributed methods. The latter requires either more GPUs and more thus communication, or gradient accumulation and thus more time. The latter also has an equal flops budget as the distributed methods. We tuned all our optimizers on the pretraining setting on a separate validation set . We also considered using SF-Nesterov, but it was hard to tune and unstable.
@@
Table 1: Optimizer hyperparameters for the three evaluated sizes. All are based on the transformer architecture, chinchilla-style [hoffmann2022chinchilla].
Results Table˜1 gives the optimal hyperparameters per scale, and Figure˜2 gives the perplexity curves. The perplexity was calculated on the C4 validation set. Consistent with the predictions of our theory, we found that an outer learning rate greater than $1.0$ performed best for SF-SGD and a relatively large effective outer learning rate also performed best for Nesterov; Moreover, acceleration consistently improved performance relative to the baseline Local SGD. In the supplementary material, we report the effect of varying the number of local steps (Section˜A.1.2), the number of clients/replicas and different ways of FLOPs allocation (Section˜A.1.3), and gradient variance (Section˜A.1.6). We also include the validation results for all the main experiments we ran in Tables˜3, 4 and 5.
@@ A.1 Language model pretraining
We study the impact of using various outer optimizers on large language model pretraining. We utilized Chinchilla-style decoder transformer architectures [hoffmann2022chinchilla] trained on the C4 dataset [c4], consistent with common practices in large-scale model training [douillard23_diloc]. The following subsections detail the specific hyperparameters, variations in training configurations (such as the number of inner steps and replicas/clients), and analyses of optimizer behavior, including learning rate scheduling and observed gradient cosine similarities.
@@ A.1.1 Hyperparameters details
We show in Table 1 the hyperparameters considered and kept, and in Table 2 the architectural hyperparameters. We use the SentencePiece tokenizer with a sequence length of $1024$ for all models. We tuned all our optimizers on a separate validation set. We also considered using the Schedule-Free Optimizer with Nesterov acceleration on top but it was hard to tune and unstable. We include the validation results for all the main experiments we ran in Tables˜3, 4 and 5.
Table 2: Model Configuration for the three evaluated sizes. All are based on the transformer architecture, chinchilla-style [hoffmann2022chinchilla].
@@ A.1.4 Schedule-free but not tuning-free
The schedule-free method of [defazio24_road_less_sched] enables not doing any learning rate scheduling, greatly simplifying training configuration. However, it doesn’t mean it is hyperparameters-tuning-free. Indeed, we found out that we had to extensively tune the initial learning rate (to $2.0$), remove learning rate warm-up contrarily to what is advised, and use a particularly low $b1$ decay: $0.2$, as illustrated in Figure 7.
@@ A.1.5 Pretraining: outer learning rate scheduling
Schedule-free SGD enables not having to manually scheduling the outer learning rate. Therefore, we wondered if we could improve the SotA federated learning baseline, DiLoCo (Nesterov outer optimizer), with an outer learning rate schedule. We investigate in Figure 7 three schedules: constant as in [douillard23_diloc], cosine decay, and linear after a plateau. For the latter we consider a constant plateau for 10% and 25% of the total steps. For each method, we also tuned the peak outer learning rate. We don’t use any warm-up in the outer optimization as we always found it to be harmful.
We find that constant outer learning rate is the best performing schedule. It’s unclear how the other schedules are interacting with the inner learning rate scheduling. A possible solution, not investigated in this report, would be to increase the number of inner steps $H$ as the inner learning rate decreases [gu2024schedulinginnersteps].
@@ A.1.6 Cosine similarity between outer gradients
Table 3: Complete hyperparameter sweep results across model scales and configurations. All experiments use C4 validation set with sequence length 1024 and batch size 512.
$H$ $M$ Algorithm Learning Rate Perplexity Model Size
Data-Parallel Baselines
- 1 Data-Parallel - 18.07 150M
- 1 Data-Parallel 4x BS 16.89 150M
- 1 Data-Parallel - 15.28 400M
- 1 Data-Parallel 4x BS 13.21 400M
- 1 Data-Parallel - 13.38 1B
- 1 Data-Parallel 4x BS 11.34 1B
Local SGD Experiments
50 4 SGD 1.0 17.75 150M
50 4 Nesterov 0.7 17.25 150M
50 4 Nesterov 1.0 16.38 150M
50 4 SF-SGD 2.0 ($\beta$=0.2) 16.88 150M
50 4 SGD 1.0 14.90 400M
50 4 Nesterov 0.7 13.71 400M
50 4 Nesterov 1.0 >>30 400M
50 4 SF-SGD 2.0 ($\beta$=0.2) 13.95 400M
50 4 SGD 1.0 13.67 1B
50 4 Nesterov 0.7 12.51 1B
50 4 SF-SGD 2.0 ($\beta$=0.2) 12.40 1B
Varying $H$ (Local Steps) at 150M, $M=4$
150 4 SGD 1.0 17.58 150M
150 4 Nesterov 0.7 17.90 150M
150 4 Nesterov 1.0 16.79 150M
150 4 SF-SGD 2.0 ($\beta$=0.2) 16.96 150M
250 4 SGD 1.0 18.20 150M
250 4 Nesterov 0.7 18.09 150M
250 4 Nesterov 1.0 17.12 150M
250 4 SF-SGD 2.0 ($\beta$=0.2) 16.97 150M
500 4 SGD 1.0 18.44 150M
500 4 Nesterov 0.7 17.95 150M
500 4 Nesterov 1.0 18.15 150M
500 4 SF-SGD 2.0 ($\beta$=0.2) 17.18 150M
1000 4 SGD 1.0 18.18 150M
1000 4 Nesterov 0.7 18.16 150M
1000 4 Nesterov 1.0 18.75 150M
1000 4 SF-SGD 2.0 ($\beta$=0.2) 17.29 150M
2000 4 SGD 1.0 18.11 150M
2000 4 Nesterov 0.7 18.40 150M
2000 4 Nesterov 1.0 18.36 150M
2000 4 SF-SGD 2.0 ($\beta$=0.2) 17.59 150M
Varying $M$ (Number of Nodes) at 150M, $H=50$
50 2 SGD 1.0 18.64 150M
50 2 Nesterov 1.0 16.81 150M
50 2 SF-SGD 2.0 ($\beta$=0.2) 17.13 150M
50 8 SGD 1.0 18.38 150M
50 8 Nesterov 1.0 16.27 150M
50 8 SF-SGD 2.0 ($\beta$=0.2) 16.92 150M
50 16 SGD 1.0 19.86 150M
50 16 Nesterov 1.0 16.25 150M
50 16 SF-SGD 2.0 ($\beta$=0.2) 16.75 150M
Table 4: Additional outer learning rate sweeps for different outer optimizers. All experiments at 150M model size with $H=50$ and $M=4$.
Algorithm Learning Rate Perplexity
SF-SGD Learning Rate Sweep ($\beta=0.2$)
SF-SGD 0.1 >>30
SF-SGD 0.5 22.89
SF-SGD 1.0 19.42
SF-SGD 1.5 18.32
SF-SGD 2.0 17.98
SF-SGD 3.0 17.96
SF-SGD 4.0 18.09
SF-SGD 5.0 17.51
Nesterov Learning Rate Sweep (Cosine Schedule)
Nesterov 0.3 17.16
Nesterov 0.5 17.06
Nesterov 0.7 16.93
Nesterov 0.9 17.19
Nesterov 1.1 17.56
SGD Learning Rate Sweep
SGD 0.3 (fixed) 21.04
SGD 0.3 (cosine) 17.68
SGD 0.5 (cosine) 16.63
SGD 0.7 (cosine) 18.84
SGD 1.0 (cosine) 19.21
Table 5: SF-SGD $\beta$ parameter sweep at 150M model size with $H=50$, $M=4$, and outer learning rate $\gamma=2.0$.
$\beta$ Value Perplexity
0.0 >>30
0.05 16.88
0.1 16.78
0.2 16.89
0.4 17.15
0.5 17.35
0.7 17.93
0.9 19.07
0.95 19.65
0.99 20.51